Learning curve problem:

A company was ordered to make 6 custom made parts. the labor costs of manufacturing all of the parts was $1000, and the manager's estimate an improvement coefficient of .6. what would be the labor costs of the next (identical) 3 part order.
Answer #1

Given That,

Improvement Coefficient = 0.6

Learning Rate = 1-0.6 = 0.4

New Labor cost for identical 3 part order = Labor Cost for 6 part order*(Learning Rate)^(ln3/ln2)

New Labor cost for identical 3 part order = 1000*(0.4)^(ln3/ln2)

New Labor cost for identical 3 part order = 1000*0.2340346 = 234.03